OpenStudy (anonymous):

Find the domain of f(x) = (all of this under a radical) 2x^2/x^2-2

OpenStudy (anonymous):

Is this it? \[\sqrt{\frac{ 2x^{2} }{x^{2}-2 }}\]

OpenStudy (anonymous):

yes!

OpenStudy (shamil98):

Set the denom = 0 , and solve for x. your domain will be all real numbers except that number.

OpenStudy (shamil98):

so \[x^2 - 2= 0\]

OpenStudy (anonymous):

This is where you should start: \[\frac{ 2x^{2} }{ x^{2}-2 }\ge0\]

OpenStudy (anonymous):

okay so I get that the bottom one would be \[x \ge2\] but what would the top be?

OpenStudy (shamil98):

I don't know what jorea is doing.. but to find the domain of that function, all you do is set the denominator inside the radical to equal zero... because dividing by zero is undefined...

OpenStudy (anonymous):

top: \[2x^{2}\ge0\] bottom:\[x^{2}-2>0\]

OpenStudy (anonymous):

i mean \[x \ge \sqrt{2}\]

OpenStudy (anonymous):

My bad on my first post.

OpenStudy (anonymous):

so where do you go from there

OpenStudy (shamil98):

\[x^2 -2 =0\] \[x = \pm \sqrt2\] so, your domain is all real numbers except positive or negative radical 2.

OpenStudy (anonymous):

oh I see thank you!